OpenStudy (anonymous):

Solve the equation. Write the smaller value of x. {x^2} - 4x - 5 = 0

OpenStudy (lynfran):

try factoring out x^2-5x+x-5=0

OpenStudy (anonymous):

x^2*x=-1.25

OpenStudy (anonymous):

it is x=-1 @Aliypop

OpenStudy (lynfran):

x(x-5)+1(x-5)=0 (x+1)=0 and (x-5)=0 now solve for x

OpenStudy (anonymous):

Don't forget to medal and fan if you already haven't. Plus send me a message if you need more help. @Aliypop

OpenStudy (lynfran):

you would get 2 values for x...

OpenStudy (lynfran):

so x=-1 and x=5 the smaller value is x=-1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry I was talking to my mom

OpenStudy (lynfran):

do you understand?

OpenStudy (anonymous):

Find the value of x. x=-20/2(-5)

OpenStudy (lynfran):

you just need to simplify to find x....what do u get?

OpenStudy (anonymous):

I can't seem to get answer for simplifying it

OpenStudy (lynfran):